# Detuning the Floquet anomalous chiral spin liquid

### v1

@@ Abstract
At high-frequency a periodically-driven quantum spin-1/2 system can emulate a chiral spin liquid (CSL) described by an effective static local chiral hamiltonian. In contrast, at low-frequency anomalous CSL can be realized in Swap Models, in which one-way spin transport occurs at the edge although the bulk time-evolution operator over one period is trivial. In this work we explicitly construct a family of Floquet quantum spin-1/2 models on the square lattice implementing Swap Models to investigate the stability of the anomalous CSL under frequency detuning and the transition to the high-frequency regime. We have used the average-energy spectrum on finite-size torus and cylinders to unfold the Floquet quasi-energy spectrum over the whole frequency range and obtain the geometrical Berry phases. This enabled us to identify three regimes upon increasing detuning: i) a finite-size regime (with no folding of the Floquet spectrum), ii) an intermediate (narrow) regime with folding and very few resonances and iii) a regime with an increased density of resonances suggesting heating. At small detuning, edge modes are revealed by spectroscopic tools and from the diamagnetic response of the system giving access to the anomalous winding number. The analysis of all the data suggests that the anomalous CSL is not continuously connected to the high-frequency CSL. We also discuss the possible occurrence of a long-lived prethermal anomalous CSL.
@@ I Introduction
Another outstanding question in the field concerns whether one can interpolate between the high-frequency and low-frequency regimes of driven spin systems within a unified framework, and, crucially, elucidate the nature of the transition between them.
Outline – In Section II we first review properties of Swap Models [8, 9] on a two-dimensional (2D) square lattice. We explicitly derive an infinite discrete set of periodic drives of XXZ hamiltonians realizing Swap Models for specific tuned values of the drive frequency. On cylinder geometries, breaking of discrete-time invariance in the stroboscopic time evolution of edge states is outlined, showing similarities with Time Crystals [10]. The investigation of the effects of detuning the frequency is carried out in Section III. A careful analysis of Floquet quasi-energy spectra and average-energy spectra on torus and cylinder geometries reveals different regimes upon increasing detuning: (i) a finite-size regime with unfolded Floquet spectrum and no energy absorption; (ii) an intermediate regime with folded Floquet spectrum but only few resonances which may signal the occurrence of a prethermal regime [11, 12], even in the thermodynamic limit; (iii) a regime where the proliferation of resonances suggests heating. In such an analysis, the average-energy concept [13], by providing a mathematical (but still physical) tool to unwind the Floquet spectrum, and the associated geometrical Berry phases, have been very helpful to characterize the various regimes. When a clear Floquet low-energy “ground-state” could be identified, signature of edge modes and anomalous spectral flow has been observed. In Section IV of this work we introduce a controlled detuning of the Swap Model, which smoothly connects to a dynamical chiral spin liquid phase in the high-frequency limit [14, 15]. Remarkably, the transition between these regimes is found to be “chaotic”, marked by the breakdown of topological protection and the disappearance of chiral edge modes due to the proliferation of resonances. This behavior stands in stark contrast to the non-interacting lattice case [2], where the edge states persist across the frequency domain under well-defined conditions.
@@ III.1 Quasi-energy spectrum and effective Floquet Hamiltonian
As a first step we use numerical calculations on finite size systems to first obtain a heuristic understanding of the role of detuning. We assume a detuned frequency $\omega=\omega_{s}+\delta\omega$ around $\omega_{s}=\omega_{0}$, keeping $\Delta=\Delta_{s}=1$. The Floquet many-body spectrum is obtained from the exact diagonalization of the Floquet unitary operator $U_{F}$. By identifying $U_{F}[t_{0}]$ with $\exp{(-iH_{F}[t_{0}]T)}$ we extract the spectrum of the Floquet hamiltonian $H_{F}[t_{0}]$ (independent of $t_{0}$) defined modulus $\omega=2\pi/T$. We show in the top panel of figure 5 the evolution within the Floquet-Brillouin zone (FBZ) of the energy spectrum on a $4\times 4$ torus, as a function of the detuning frequency. As seen the bandwidth increases linearly with $\delta\omega$. This defines a finite-size cross-over detuning frequency $\delta\omega_{\rm fold}(N)\propto J/N$ above which the winding of the Floquet spectrum in the quasi-energy FBZ occurs. For $N=16$ sites we find that $\delta\omega_{\rm fold}\simeq 0.05$. The frequency interval $\omega\in[0.45,0.56]$ can then be considered as a “finite-size” regime characterized by a limited extension of the Floquet spectrum within the quasi-energy FBZ.
@@
Note that $H_{F}^{\Delta}(\omega)$ is (in lowest order) invariant under time reversal symmetry (i.e it is not chiral) but is odd in $\delta\omega$. The expansion at first order contains interactions at length $1$ (bonds $a$ and $d$) and $\sqrt{5}$ (bonds $b^{\prime}$ and $c^{\prime}$) as shown in figure 6. This pattern has the same spatial symmetries as the drive Hamiltonian: translational symmetries of the square lattice are partially broken, the system being invariant only under $T_{x}^{2}$, $T_{y}^{2}$ and $T_{x}\,T_{y}=T_{y}\,T_{x}$. Also the $\pi/2$-rotation is broken. Note that the Floquet hamiltonian $H_{F}[t]$ depends on the initial time $t$. E.g. $H_{F}[T/4]$ is the same as $H_{F}\equiv H_{F}[0]$ shown in figure 6 (left panel) but rotated by $\pi/2$ around a lattice site.
@@
Note that, on the cylinder, the 1st-order bulk (non-chiral) local Floquet Hamiltonian is not able to reproduce the chiral Floquet states. This shows that the Floquet spectrum on a system with boundaries cannot be obtained from an expansion of the bulk Floquet Hamiltonian alone, but an additional non-local edge Hamiltonian is necessary. This is to be contrasted with static and local chiral Heisenberg models whose many-body spectra show edge modes on geometries with edges, like a cylinder [18, 19], a disk [20] or a half-infinite plane [21]. In this case the chiral modes appear within the many-body bulk energy gap and no sub-band structure like in figure 7 is present.
@@ III.2 Average-energy spectrum
We shall first concentrate on small detuning, typically within the finite-size frequency interval $|\delta\omega|<\delta\omega_{\rm fold}(N)$ or just slightly beyond. In that regime, the average-energy spectrum computed on a $4\times 4$ torus is shown on the middle panel of figure 5 and compared to the Floquet quasi-energy spectrum on the same system. Interestingly we find that the average-energy bandwidth is always of order $4.5J$ ($\sim 9\omega$) even in the limit $\delta\omega\rightarrow 0$, in stark contrast to the Floquet spectrum’s bandwidth which scales as $\delta\omega$. In fact, at small $\delta\omega$, we observe a close similarity between the average-energy spectrum $\{{\text{\ae }}_{n}\}$ and the (rescaled) Floquet spectrum $\{{\cal E}_{n}\}$. This suggests a simple relation between these spectra at leading order in $\delta\omega$,
@@
The average-energy spectrum on a $4\times 4$ cylinder shown on the middle panel of figure 7 bears similar features as the torus case. However, the low-energy part of the spectrum shows many more closely packed levels which may be associated to edge states. Also, the lowest average-energy states (the ferromagnetic state) are (is) located at the bottom (top) or at the top (bottom) of the Floquet subbands in the finite-size regime for $\delta\omega<0$ or $\delta\omega>0$, respectively. All levels evolve smoothly and continuously with increasing detuning and, surprisingly, even beyond the finite-size regime. We have checked numerically that, in contrast to the torus geometry, the average-energy spectrum cannot be described by the Hamiltonian $\frac{J}{4}{\hat{h}}_{\sqrt{5}}$ in the small $\delta\omega$ limit. In particular, it is easy to see that the exact average-energy (2NL)J/16(2N-L)J/16 of the fully polarized ferromagnetic state (an exact Floquet eigenstate) is different from its energy $(2N-\frac{3}{2}L)J/16$ in the spectrum of $\frac{J}{4}{\hat{h}}_{\sqrt{5}}$ (upper bounds in both spectra). However a simple relation between the average-energy spectrum and the Floquet spectrum holds in lowest order in $\delta\omega$ (see Appendix B.4):
@@
where, in contrast the the torus case, there is an energy shift ${\epsilon_{P}}=-P\frac{\omega}{2}$ to move the $P$ Floquet sub-bands (here the $P=1$ sub-band) down to zero (average) energy, overlapping with the $P=0$ sub-band. Note also that the $H_{abcd}$ Hamiltonian entering in the expression of the average-energy spectrum and in $H_{F}(\omega)$ (see Appendix B) now involves, in addition to the bulk part $\frac{J}{4}{\hat{h}}_{\sqrt{5}}$, long-range terms acting on the boundaries.
@@
whose behavior consistently supports the existence of three regimes. The geometric phases are shown in the bottom panels of figure 5 and 8 on the same $4\times 4$ torus and figure 7 and 9 on the cylinder and for the same selection of low average-energy states. While the (highly excited) ferromagnetic state has zero Berry phase (since $T{\cal E}_{\rm ferro}=T{\text{\ae }}_{\rm ferro}\hskip 2.84544pt{\rm mod}[2\pi]$) the geometrical phases of the low average-energy states are finite, even in the $\delta\omega\rightarrow 0$ limit. The latter tend to wind by $2\pi$ when moving away from the fine tuned frequency $\omega=\omega_{s}$. We interpret these features as being signatures of the anomalous character of the CSL around $\omega=\omega_{s}$. However, the occurrence of more frequent resonances when $|\delta\omega|$ increases above $\delta\omega_{\times}$ leads to an increasingly erratic behavior of the geometric phases.
@@ III.4 Anomalous winding number
Figure 12: $U_{F}$ spectra of a $4\times 4$ cylinder in the complex plane for $\omega=0.5$ (left panels) and $\omega=0.53$ (right panels) as a function of $\theta=2\pi\phi/N_{\text{pl}}$, for a total flux $\phi$ varying from 0 (centers of plots) to $N_{\rm pl}=12$ (top) or $N_{\rm pl}/2=6$ (bottom) flux quanta (perimeters of plots). The top panels correspond to $S_{z}=N/2-1$ i.e. a single spin-flip described by the tight-binding model of [3]. The bottom panels correspond to the $S_{z}=0$ sector describing a half-filled interacting hard-core boson system. Although all eigenvalues have a modulus 1, we represent them on radius-$\phi$ concentric circles to give a visual representation of the states flowing through the various chosen branch-cuts (represented as red lines at angles $\Phi_{\text{cut}}=\pi,\pi/2,7\pi/8,-7\pi/8$). Values of $\theta$ multiple of $\pi/6$ corresponding to integer values of flux quanta threading the whole sample are indicated.
As a simple test case, we shall first consider the case of a single spin flip ($S_{z}=N/2-1$) whose dynamics is described by the (non-interacting) single particle tight-binding model introduced in [3]. The diamagnetic response of this model has been studied in [28] for an open geometry in the plane. We show here that similar results are obtained in our setting on the cylinder. Two typical spectra of $U_{F}$ are shown in the complex plane in the top panels of figure 12 at the fine-tuned frequency and for a small detuning of $6\%$. At fine-tuning, we observe $N_{\rm pl}=12$ degenerate bulk states (localized on plaquettes) of eigenvalue $\exp{(-i\theta)}$ and $L_{x}/2=2$ edge modes on both left (L) and right (R) edges. As expected the L edge modes (of eigenvalues $\pm 1$) do not show any flux dependence while the two R edge modes have eigenvalues $\pm\exp{(6i\theta)}$. In our representation of figure 12 the bulk states and the R edge modes give two sets of spirals winding around the center of the plots in opposite directions (and with different pitch angles). Note that the R edge states switch their eigenvalues upon increasing the flux $\phi$ by one unit (i.e. $\theta$ by $\pi/6$). The edge state contribution to $N_{1}$ in Eq. (33) is therefore a periodic function of the flux of period $\Delta\phi_{\rm edge}=1$, as also found in [28] for a planar geometry. In contrast, the contribution from the bulk states is periodic of period $\Delta\phi_{\rm bulk}=N_{\rm pl}=12$. When a small detuning is introduced we find that the exact degeneracy of the bulk states is lifted and that bulk and edge modes get coupled.
@@
Next, we consider the $S_{z}=0$ sector, as in other parts of this work, to compute the trace in (33). In contrast to the single particle case, we have now a half-filled system of hard-core bosons and the nearest-neighbor coupling $S_{i}^{z}S_{j}^{z}$ (corresponding to a repulsion between two neighboring bosons) becomes relevant and plays a key role, even at the fine-tuned frequency $\omega_{s}=0.5$: edge and bulk states are coupled by an arbitrary small flux and the tensor product structure (14) of $U_{F}(\omega_{s})$ breaks down.
@@
Figure 13: Winding number $N_{1}(\phi)$ versus $\phi$ for the single spin-flip case (two top panels) and the $S_{z}=0$ sector (two bottom panels), in the flux interval $[0,12]$ and $[0,6]$, respectively, and for $\omega=0.5$ and $\omega=0.53$. Data obtained for the four different branch-cuts of figure 12 are shown using different colors as indicated on the bottom plot. To eliminate a trivial dependence of $N_{1}(0)$ with the choice of the branch-cut, all data have been shifted by $N_{1}(0)$ for clarity. Then, all data points displayed have integer values, even away from fine-tuning.
@@
Results for the spectra of the unitary operator $U_{F}(\omega,\phi)$ versus the flux $\phi$ are shown in the bottom row of figure 12, for the same frequencies $\omega=0.5$ and $0.53$ as for the single spin-flip case. We restrict here to $\phi\in[0,6]$ since the spectra are symmetric under $\phi\rightarrow N_{\rm pl}-\phi$ ($N_{\rm pl}=12$). Since we have a many-body system whose Hilbert space size scales exponentially with the system size, the density of points is much larger than in the one-particle case, which makes it now impossible to follow by eye the individual state trajectories. At $\omega=0.5$ numerous gaps (the white regions) are observed in the spectrum (left panel). The two white blobs at the center of the figure corresponds to the $\phi=0$ gap (separating the two degenerate sub-bands at quasi-energies 0 and $\pm\omega/2$) which survives up to a small flux value $\phi\sim 0.45$. The gaps appearing at $\phi>1$ are due to lattice commensurability effects, like in the Hofstadter tight-binding model, made here relevant by the fact that (most) bulk states become spatially extended, in contrast to the single particle case. Both types of gaps are rapidly reduced/washed out by a small detuning, as seen on the right panel. Note that we observe that a small number (=346=346) of states are flux independent with $U_{F}$ eigenvalues $\pm 1$ (i.e. on the horizontal axis of figure 12). We believe that they correspond to a (small) subset of the edge states strictly localized on the left boundary.
@@
As for the single particle case, we have computed $N_{1}(\phi)$ from the trace formula (33) for the same choices of the branch-cuts and for the same frequencies $\omega=0.5$ and $\omega=0.53$. Note that, when the branch-cut is chosen as the negative horizontal axis (continuous red line), we have now to pay attention to the fact that some states are located exactly on this axis (for $\omega=0.5$). Our “traditional” choice of the first FBZ $\omega/2,\omega/2]$ corresponds in fact to the branch-cut lying just below the negative horizontal axis. A very close look of the results shown in figure 13 reveals that $N_{1}$ versus $\phi$ has a staircase behavior as expected ($N_{1}$ is quantized for $\omega=0.5$ as for the single particle case), although with such a high density of integer jumps that it looks like a continuous function. We also see that the choice of the branch-cut matters quantitatively. However, we observe that $N_{1}(\phi)$ has generically some flat portions in the center of the plots, especially for the two branch cuts corresponding to the horizontal and vertical semi-infinite axis. This suggests that $W_{A}=0$ and we argue that this is expected for a half-filled hardcore bosonic systems: A particle-hole transformation in the bosonic language corresponds to a $S_{z}\rightarrow-S_{z}$ transformation in the spin language. As explicitly seen above in the case of a single spin-flip, a particle-hole transformation leads to $W_{A}\rightarrow-W_{A}$. Since, the $S_{z}=0$ sector is particle-hole symmetric, we then expect $W_{A}=0$. Finally we note that under detuning the flat regions of $N_{1}(\phi)$ are quickly altered.

### v2

@@ Abstract
At high-frequency a periodically-driven quantum spin-1/2 system can emulate a chiral spin liquid (CSL) described by an effective static local chiral Hamiltonian. In contrast, at low-frequency these settings realize ”Swap” models exhibiting anomalous CSL phases, in which one-way spin transport occurs at the edge although the bulk time-evolution operator over one period is trivial. In this work we explicitly construct a family of Floquet quantum spin-1/2 models on the square lattice implementing Swap models to investigate the stability of the anomalous CSL under frequency detuning and the transition to the high-frequency regime. We used the average-energy spectrum on finite-size torus and cylinders to unfold the Floquet quasi-energy spectrum over the whole frequency range and obtain the geometrical Berry phases. This enabled us to identify three regimes upon increasing detuning: i) a finite-size regime (with no folding of the Floquet spectrum), ii) an intermediate (narrow) regime with folding and very few resonances and iii) a regime with an increased density of resonances suggesting heating. At small detuning, edge modes are revealed by spectroscopic tools and from the Streda response of the system giving access to the anomalous winding number. The analysis of all the data suggests that the anomalous CSL is not continuously connected to the high-frequency CSL. We also discuss the possible occurrence of a long-lived prethermal anomalous CSL.
@@ I Introduction
Another outstanding issue in the field concerns whether one can interpolate between the high-frequency and low-frequency regimes of driven spin systems within a unified framework, and, crucially, elucidate the nature of the transition between them.
In this work, in order to address the two fundamental issues described above, we have adopted the following outline. In Section II we first review properties of Swap models [8, 9] on a two-dimensional (2D) square lattice. We explicitly derive an infinite discrete set of periodic drives of XXZ Hamiltonians realizing Swap models for specific tuned values of the drive frequency. On cylinder geometries, breaking of discrete-time invariance in the stroboscopic time evolution of edge states is outlined, showing similarities with Time Crystals [10]. The investigation of the effects of detuning the frequency is carried out in Section III. A careful analysis of Floquet quasi-energy spectra and average-energy spectra on torus and cylinder geometries reveals different regimes upon increasing detuning: (i) a finite-size regime with unfolded Floquet spectrum and no energy absorption; (ii) an intermediate regime with folded Floquet spectrum but only few resonances which may signal the occurrence of a prethermal regime [11, 12], even in the thermodynamic limit; (iii) a regime where the proliferation of resonances suggests heating. In such an analysis, the average-energy concept [13], by providing a mathematical (but still physical) tool to unwind the Floquet spectrum, and the associated geometrical Berry phases, have been very helpful to characterize the various regimes. When a clear Floquet low-energy “ground-state” could be identified, signature of edge modes and anomalous spectral flow has been observed. In Section IV of this work we introduce a controlled detuning of the Swap model, which smoothly connects to a dynamical chiral spin liquid phase in the high-frequency limit [14, 15]. Remarkably, the transition between these regimes is found to be “chaotic”, marked by the breakdown of topological protection and the disappearance of chiral edge modes due to the proliferation of resonances. This behavior stands in stark contrast to the non-interacting lattice case [2], where the edge states persist across the frequency domain under well-defined conditions.
@@ III.1 Quasi-energy spectrum and effective Floquet Hamiltonian
As a first step we use numerical calculations on finite size systems to obtain a heuristic understanding of the role of detuning. We assume a detuned frequency $\omega=\omega_{s}+\delta\omega$ around $\omega_{s}=\omega_{0}$, keeping $\Delta=\Delta_{s}=1$. The Floquet many-body spectrum is obtained from the exact diagonalization of the Floquet unitary operator $U_{F}$. By identifying $U_{F}[t_{0}]$ with $\exp{(-iH_{F}[t_{0}]T)}$ we extract the spectrum of the Floquet hamiltonian $H_{F}[t_{0}]$ (independent of $t_{0}$) defined modulus $\omega=2\pi/T$. We show in the top panel of figure 5 the evolution within the Floquet-Brillouin zone (FBZ) of the energy spectrum on a $4\times 4$ torus, as a function of the detuning frequency. As seen the bandwidth increases linearly with $\delta\omega$. This defines a finite-size cross-over detuning frequency $\delta\omega_{\rm fold}(N)\propto J/N$ above which the winding of the Floquet spectrum in the quasi-energy FBZ occurs. For $N=16$ sites we find that $\delta\omega_{\rm fold}\simeq 0.05$. The frequency interval $\omega\in[0.45,0.56]$ can then be considered as a “finite-size” regime characterized by a limited extension of the Floquet spectrum within the quasi-energy FBZ.
@@
Note that $H_{F}^{\Delta}(\omega)$ is (in lowest order) invariant under time reversal symmetry (i.e it is not chiral) but is odd in $\delta\omega$. The expansion at first order contains interactions at length $1$ (bonds $a$ and $d$) and $\sqrt{5}$ (bonds $b^{\prime}$ and $c^{\prime}$) as shown in figure 6. This pattern has the same spatial symmetries as the drive Hamiltonian: translational symmetries of the square lattice are partially broken, the system being invariant only under $T_{x}^{2}$, $T_{y}^{2}$ and $T_{x}\,T_{y}=T_{y}\,T_{x}$. Also the $\pi/2$-rotation is broken. Note that the Floquet hamiltonian $H_{F}[t_{0}]$ depends on the initial time $t_{0}$. E.g. $H_{F}[T/4]$ is the same as $H_{F}\equiv H_{F}[0]$ shown in figure 6 (left panel) but rotated by $\pi/2$ around a lattice site.
@@
Note that, on the cylinder, the 1st-order bulk (non-chiral) local Floquet Hamiltonian is not able to reproduce the chiral Floquet states. This shows that the Floquet spectrum on a system with boundaries cannot be obtained from an expansion of the bulk Floquet Hamiltonian alone, but an additional non-local edge Hamiltonian is necessary. This is to be contrasted with static and local chiral Heisenberg models whose many-body spectra show edge modes on geometries with edges, like a cylinder [18, 19], a disk [20] or a half-infinite plane [21]. In this case the chiral modes appear within the many-body bulk energy gap and no sub-band structure like in figure 7 is present. Overall, this reflects the ‘Floquet anomalous’ nature of the edge modes depicted in Figure 7.
@@ III.2 Average-energy spectrum
We shall first concentrate on small detuning, typically within the finite-size frequency interval $|\delta\omega|<\delta\omega_{\rm fold}(N)$ or just slightly beyond. In that regime, the average-energy spectrum computed on a $4\times 4$ torus is shown on the middle panel of figure 5 and compared to the Floquet quasi-energy spectrum on the same system. Interestingly we find that the average-energy bandwidth remains constant (of order $4.5J\sim 9\omega$ for the parameters of Fig. 5) even in the limit $\delta\omega\rightarrow 0$, in stark contrast to the Floquet spectrum’s bandwidth which scales as $\delta\omega$. In fact, at small $\delta\omega$, we observe a close similarity between the average-energy spectrum $\{{\text{\ae }}_{n}\}$ and the (rescaled) Floquet spectrum $\{{\cal E}_{n}\}$. This suggests a simple relation between these spectra at leading order in $\delta\omega$,
@@
The average-energy spectrum on a $4\times 4$ cylinder shown on the middle panel of figure 7 bears similar features as the torus case. However, the low-energy part of the spectrum shows many more closely packed levels which may be associated to edge states, as suggested by edge mode spectroscopy (see next in III.3). Also, the lowest average-energy states (the ferromagnetic state) are (is) located at the bottom (top) or at the top (bottom) of the Floquet subbands in the finite-size regime for $\delta\omega<0$ or $\delta\omega>0$, respectively. All levels evolve smoothly and continuously with increasing detuning and, surprisingly, even beyond the finite-size regime. We have checked numerically that, in contrast to the torus geometry, the average-energy spectrum cannot be described by the Hamiltonian $\frac{J}{4}{\hat{h}}_{\sqrt{5}}$ in the small $\delta\omega$ limit. In particular, it is easy to see that the exact average-energy (2NL)J/16(2N-L)J/16 of the fully polarized ferromagnetic state (an exact Floquet eigenstate) is different from its energy $(2N-\frac{3}{2}L)J/16$ in the spectrum of $\frac{J}{4}{\hat{h}}_{\sqrt{5}}$ (upper bounds in both spectra). However a simple relation between the average-energy spectrum and the Floquet spectrum holds in lowest order in $\delta\omega$ (see Appendix B.4):
@@
where, in contrast the torus case, there is an energy shift ${\epsilon_{P}}=-P\frac{\omega}{2}$ to move the $P$ Floquet sub-bands (here the $P=1$ sub-band) down to zero (average) energy, overlapping with the $P=0$ sub-band. Note also that the $H_{abcd}$ Hamiltonian entering in the expression of the average-energy spectrum and in $H_{F}(\omega)$ (see Appendix B) now involves, in addition to the bulk part $\frac{J}{4}{\hat{h}}_{\sqrt{5}}$, long-range terms acting on the boundaries.
@@
whose behavior consistently supports the existence of three regimes. The geometric phases are shown in the bottom panels of figure 5 and 8 on the same $4\times 4$ torus and figures 7 and 9 on the cylinder and for the same selection of low average-energy states. While the (highly excited) ferromagnetic state has zero Berry phase (since $T{\cal E}_{\rm ferro}=T{\text{\ae }}_{\rm ferro}\hskip 2.84544pt{\rm mod}[2\pi]$) the geometrical phases of the low average-energy states are finite, even in the $\delta\omega\rightarrow 0$ limit. The latter tend to wind continuously around the circle when moving away from the fine tuned frequency $\omega=\omega_{s}$. We interpret these features as being signatures of the anomalous character of the CSL around $\omega=\omega_{s}$. However, the occurrence of more frequent resonances when $|\delta\omega|$ increases above $\delta\omega_{\times}$ leads to an increasingly erratic behavior of the geometric phases.
@@ III.4 Anomalous winding number
Let us also mention that, in the following, we shall consider particular $S_{z}$ sectors corresponding to given hard-core boson fillings. Therefore the trace in the formula above will be a partial trace over the corresponding reduced Hilbert space and, in that case, ${\rm Tr}[H_{\rm drive}(t)]$ no longer vanishes. However, the magnetic flux appears only in the off-diagonal elements of $H(t)$ and ${\rm Tr}[H_{\rm drive}(t)]$ is just a flux-independent constant that does not contribute to $W_{A}$.
@@
Figure 12: $U_{F}$ spectra of a $4\times 4$ cylinder in the complex plane for $\omega=0.5$ as a function of $\theta=2\pi\phi/N_{\text{pl}}$, for a total flux $\phi$ varying from 0 (central plots) to $N_{\rm pl}=12$ (left) or $N_{\rm pl}/2=6$ (right) flux quanta (perimeters of plots). The four panels, from left to right, correspond to a single spin-flip ($S_{z}=N/2-1=7$) described by the tight-binding model of [3], two ($S_{z}=N/2-2=6$) and three ($S_{z}=N/2-3=5$) spin-flips and a half-filled ($S_{z}=0$) interacting hard-core boson system, respectively. Although all eigenvalues have a modulus 1, we represent them on radius-$\phi$ concentric circles to give a visual representation of the states flowing through the various chosen branch-cuts (represented as red lines at angles $\Phi_{\text{cut}}=\pi,\pi/2,7\pi/8,-7\pi/8$). Values of $\theta$ multiple of $\pi/6$ corresponding to integer values of flux quanta threading the whole sample are indicated.
Figure 13: Same as Fig. 12 for a small detuning $\omega=0.53$.
As a simple test, we shall first consider the case of a single spin flip ($S_{z}=N/2-1$) whose dynamics is described by the (non-interacting) single particle tight-binding model introduced in [3]. The diamagnetic response of this model has been studied in [28] for an open geometry in the plane. We show here that similar results are obtained in our setting on the cylinder. Two typical spectra of $U_{F}$ are shown in the complex plane in the left panel of figure 12 at the fine-tuned frequency and in the left panel of figure 13 for a small detuning of $6\%$. At fine-tuning, we observe $N_{\rm pl}=12$ degenerate bulk states (localized on plaquettes) of eigenvalue $\exp{(-i\theta)}$ and $L_{x}/2=2$ edge modes on both left (L) and right (R) edges. As expected the L edge modes (of eigenvalues $\pm 1$) do not show any flux dependence while the two R edge modes have eigenvalues $\pm\exp{(6i\theta)}$. In our representation of figure 12 the bulk states and the R edge modes give two sets of spirals winding around the center of the plots in opposite directions (and with different pitch angles). Note that the R edge states switch their eigenvalues upon increasing the flux $\phi$ by one unit (i.e. $\theta$ by $\pi/6$). The edge state contribution to $N_{1}$ in Eq. (33) is therefore a periodic function of the flux of period $\Delta\phi_{\rm edge}=1$, as also found in [28] for a planar geometry. In contrast, the contribution from the bulk states is periodic of period $\Delta\phi_{\rm bulk}=N_{\rm pl}=12$. When a small detuning is introduced we find that the exact degeneracy of the bulk states is lifted and that bulk and edge modes get coupled.
@@
Next, we turn to the case of two and three spin flips, i.e. $S_{z}=N/2-2=6$ and $S_{z}=N/2-3=5$ respectively, which can be viewed as $N_{p}=2$ and $N_{p}=3$ interacting hard-core bosons in our finite cylinder. For completeness we also consider the $S_{z}=0$ sector, as in the previous parts of this work, which corresponds to an interacting half-filled hard-core bosonic system ($N_{p}=8$ bosons on $N=16$ sites). In contrast to the single particle case, the nearest-neighbor Ising coupling $S_{i}^{z}S_{j}^{z}$ (corresponding to a repulsion between two neighboring bosons) becomes relevant and plays a key role, even at fine-tuned frequency $\omega_{s}=0.5$ for which the tensor product structure (14) of $U_{F}(\omega_{s})$ breaks down.
@@
Figure 14: Winding number $N_{1}(\phi)$ versus $\phi$ for $N_{p}=1$ (top panels), $N_{p}=2$, $N_{p}=3$ (middle panels) and $N_{p}=N/2=8$ (bottom panels) hard-core bosons, in the flux intervals $[0,12]$ (top and middle panels) or $[0,6]$ (bottom panels), for $\omega=0.5$ (left panels) and $\omega=0.53$ (right panels). Data obtained for the four different branch-cuts of figure 12 are shown using different colors as indicated on the bottom-right plot. To eliminate a trivial dependence of $N_{1}(0)$ with the choice of the branch-cut, all data have been shifted by $N_{1}(0)$ for clarity. Then, all data points displayed have integer values, even away from fine-tuning. Estimations of the slope $W_{A}$ are shown on the plots as dashed lines. For $N_{p}=3$ the dashed line is only a guide to the eye.
@@
Results for the spectra of the many-body unitary operator $U_{F}(\omega,\phi)$ at the fine-tuned frequency $\omega=0.5$ are shown versus the flux $\phi$ in the three right panels of figure 12. For $S_{z}=0$ we restrict here to $\phi\in[0,6]$ since the spectra are symmetric under $\phi\rightarrow N_{\rm pl}-\phi$ ($N_{\rm pl}=12$). Since we have now many-body systems whose Hilbert space size scales exponentially with system size, the density of points becomes larger than in the one-particle case, especially for $S_{z}=0$ for which it is now impossible to follow by eye the individual state trajectories. However, for the case of two and three particles, one can still identify (5 and 9) degenerate edge modes forming spirals winding counterclockwise and (3 and 4) degenerate ”bulk modes” winding in the opposite direction. The degeneracy of these modes indicates some degree of localization in the bulk and/or on the boundary. For $S_{z}=0$ (half-filled bosonic system), the density of points is very high (forming a quasi-continuum) and numerous gaps (the white regions) are observed in the spectrum (right panel). The two white blobs at the center of the figure corresponds to the $\phi=0$ gap (separating the two degenerate sub-bands at quasi-energies 0 and $\pm\omega/2$) which survives up to a small flux value $\phi\sim 0.45$. The gaps appearing at $\phi>1$ are due to lattice commensurability effects, like in the Hofstadter tight-binding model, made here relevant by the fact that (most) bulk states become spatially extended, in contrast to the single particle case.
Similar spectra are shown in the three right panels of figure 13 for the same detuned frequency $\omega=0.53$ as for the single particle case (left panel). The detuning, although quite small, has a strong effect on the spectrum: the degeneracies of the edge and bulk modes are immediately lifted and edge and bulk states couple to each other. At $S_{z}=0$, gaps are rapidly reduced/washed out, as seen on the far-right panel. Note that, in this case, we observe that a small number (=346=346) of states are flux independent with $U_{F}$ eigenvalues $\pm 1$ (i.e. on the horizontal axis of figure 13). We believe that they correspond to a (small) subset of the edge states strictly localized on the left boundary.
As for the single particle case, we have computed $N_{1}(\phi)$ from the trace formula (33) for the same choices of the branch-cuts and for the same frequencies $\omega=0.5$ and $\omega=0.53$. A close look of the results shown in figure 14 reveals that $N_{1}$ versus $\phi$ has always a staircase behavior with integer jumps as expected ($N_{1}$ itself is quantized for $\omega=0.5$ for all particle number $N_{p}$). The behavior of $N_{1}(\phi)$ for $N_{p}=2$ is qualitatively similar to the single particle case with well-defined linear portions from which it is easy to extract the anomalous spectral flow $W_{A}=13$. The case $N_{p}=3$ gives a larger value around $W_{A}=88$ but a precise determination is no longer possible.
@@
The $S_{z}=0$ ($N_{p}=N/2$) case bears some special features and is qualitatively different from the previous cases. First, we see that the density of integer jumps is so high that $N_{1}(\phi)$ looks now like a continuous function. We also note that, when the branch-cut is chosen as the negative horizontal axis (continuous red line), we have now to pay attention to the fact that some states are located exactly on this axis (for $\omega=0.5$). Our “traditional” choice of the first FBZ $\omega/2,\omega/2]$ corresponds in fact to the branch-cut lying just below the negative horizontal axis. We observe that $N_{1}(\phi)$ has generically some flat portions in the center of the plots, especially for the two branch cuts corresponding to the horizontal and vertical semi-infinite axis. This suggests that $W_{A}=0$ and we argue that this is expected for a half-filled hardcore bosonic systems: A particle-hole transformation in the bosonic language corresponds to a $S_{z}\rightarrow-S_{z}$ transformation in the spin language. As explicitly seen above in the case of a single spin-flip, a particle-hole transformation leads to $W_{A}\rightarrow-W_{A}$. Since, the $S_{z}=0$ sector is particle-hole symmetric, we then expect $W_{A}=0$.
Finally, we stress that the effect of detuning is much more severe in the interacting case than for a single particle. For $N_{p}\geq 2$ the linear behaviors of $N_{1}(\phi)$ are quickly altered by detuning, even very small, as seen in the right panels of figure 14. This shows that the Streda response is a very sensitive tool to probe the stability of anomalous interacting topological phases.